Suppose that the current value of the S&P 500 stock index is USD 3000. Assume that the rates of interest in USD and Euro are respectively 3% and 1% per year with continuous compounding, and that the S&P 500 index pays a continuous dividend rate of 2% per year. Finally, the spot exchange rate is USD 1.1 per Euro.

• Compute the forward price of the S&P 500 in USD for delivery in one year.

• Compute the forward price (exchange rate) of USD in Euro for delivery in one year.

• Compute the forward price of the S&P 500 in Euro for delivery in one year.
Answer #1

Answer a) Forward price of  the S&P 500 in USD = S*ert = 3000* e0.03*1 = $ 3091.42

Answer b) Forward price (exchange rate) of USD in Euro = S *(1 + Id) / (1 + Ie)

F= 1.1*(1.03/1.01) = 1.1218.

Answer c) Forward price of  the S&P 500 in Euro = $ 3091.42 / 1.1218= Euro 2755.813